Question Part Points Submissions Used 1 2 3 0/1 –/1 –/1 1/15 0/15 0/15 Total 0/3 Use the Trapezoidal Rule, the Midpoint Rule, and Simpson's Rule to approximate the given integral with the specified value of n. (Round your answers to six decimal places.) 4 e2√t sin(4t) dt, n = 8
Answer #1

Qua TAShezoila Rule.- AN- ket (1-472 hore h h Thesefore At = 4.0 = ! Divide hs bihtcavd wi ow chA 3 aso Gunchin Now, we ELt.Mid hoiht Raule: Cot ) It stalc kat uy b-a At Hse ao, b- 4, h-s 4-0 ! Dvie Intoono Le,] nto wik enl hit Nou w (3= in (2) - 0.imoh Rue:- 914le 1Cat 4-0 Div ne h htesv& azo NO, 4-46067 1eu, 96 1 3 6 ) = -1.945 24 7 5151407 4 (2)4 4E3 (?) - 167- oi to13